# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ V Conclusion
Remote sensing image interpretation technology is undergoing a paradigm shift from ”vision-centered perception” to ”language-centered cognition.” This paper traces the dialectical relationship between language and vision in the human cognitive system, revealing the inherent limitations of the current vision-centered paradigm in terms of semantic abstraction, dynamic feedback, and closed-loop cognition. Grounded in the Global Workspace Theory of cognitive science, we propose a novel interpretation paradigm that positions language as the central hub for multimodal representation. This framework not only reconstructs the full perception–cognition–decision loop, but also empowers interpretation systems with cross-modal semantic association and interpretable reasoning through the referential and logical predicate functions of language. Compared to traditional vision-centered approaches, the language-centered framework presents three transformative advantages: The natural language interface for task understanding eliminates barriers in human-machine interaction. The integration of dynamic knowledge graphs overcomes the limitations of perception constrained by ”what is seen is what is known.” The cyclic interaction mechanism of the cognitive workspace enables a qualitative leap from static interpretation to autonomous decision-making. While current technologies still face challenges in multimodal representation alignment and spatiotemporal knowledge evolution modeling, the advancement of a language-centered cyclic interpretation paradigm holds promise for transcending the instrumental limitations of existing systems, steering remote sensing interpretation toward becoming a cognitive agent capable of self-evolution. Future research should focus on deep coupling between language, vision, and physical space, constructing dynamically evolving cognitive schemata under spatiotemporal constraints. Ultimately, this will enable a leap from data-driven to emergent cognitive intelligence in remote sensing, opening a new paradigm of cognitive computing where ”language reconstructs the world.”

### v2

@@ VI Conclusion
Grounded in the GWT of cognitive science, we propose a novel interpretation paradigm that positions language as the central hub for multimodal representation. This framework not only reconstructs the full perception–cognition–decision loop but also empowers interpretation systems with cross-modal semantic association and interpretable reasoning through the referential and logical predicate functions of language. Compared with the visual-centered paradigm, the language-centered framework presents three transformative advantages: The natural language interface for task understanding eliminates barriers in human-machine interaction. The integration of dynamic knowledge graphs overcomes the limitations of perception constrained by “what is seen is what is known.” The cyclic interaction mechanism of the cognitive workspace enables a qualitative leap from static interpretation to autonomous decision-making. While current technologies still face challenges in multimodal representation alignment and spatiotemporal knowledge evolution modeling, the advancement of a language-centered cyclic interpretation paradigm holds promise for overcoming the instrumental limitations of existing systems, steering remote sensing interpretation toward becoming a cognitive agent capable of self-evolution. Future research should focus on deep coupling between language, vision, and physical space, constructing dynamically evolving cognitive schemata under spatiotemporal constraints. Ultimately, this effort will enable a leap from data-driven to emergent cognitive intelligence in remote sensing, creating a new paradigm of cognitive computing where “language reconstructs the world.”